OpenStudy (anonymous):

Convert to a decimal: 90 2/5%

OpenStudy (ahsome):

Is it: \[90\frac{2}{5}\] ?

OpenStudy (ahsome):

Or: \[\frac{902}{5}\]?

OpenStudy (ahsome):

Hello, @jillarie8 ?

OpenStudy (ahsome):

Lets assume you are asking how to conver \(90.5\) to decimal Thats easy, simply divide by 100 \[Decimal=\frac{90.5}{100}\]\[Decimal=0.905\]

OpenStudy (ahsome):

The decimal equalts the percentage divided by 100 \[Decimal=\frac{Percentage}{100}\]

OpenStudy (ahsome):

\[Decimal=\frac{90\frac{2}{5}}{100}\]\[Decimal=0.905\]

OpenStudy (shinalcantara):

i presume that it is a mixed number \[90 \frac{ 2 }{ 5 } \] Changing the one in the fraction form into decimal you'll have: 2/5 = 0.4 Then you'll have it as 90.4% To express it into decimal form, move two decimal places to the left 90.4% = 0.904

OpenStudy (ahsome):

Thats right @shinalcantara :)